R^3 has infinitely many non zero, proper vector subspaces. True or false with full explanation



1
Expert's answer
2021-12-09T13:45:51-0500

Let "\left\{e_{1}, e_{2}\right\}" be a basis in "\mathbb{R}^{3}" . Then

"W_{n}=\left\{\alpha\left(e_{1}+n e_{2}\right), \alpha \in \mathbb{R}\right\}, n \in \mathbb{N}"

are all different proper vector subspaces. Indeed, suppose "W_{n}=W_{m}" .

Then "e_{1}+n e_{2}=\alpha\left(e_{1}+m e_{2}\right)" for some "\alpha."

Then "(\alpha-1) e_{1}+(m \alpha-n) e_{2}=0"

From which "\alpha=1, \alpha=n / m, i.e. \ n=m."

This proves that "W_{n}" are all different. And there are infinitely many of them.

Conclusion: "\mathbb{R}^{3}" has infinitely many nonzero, proper vector subspaces - true.